A train started from rest and moved with constant acceleration.at one time it was traveling 30m/s and 160m farther on it was traveling 50m/s.
Calculate
1.the acceleration
2.the time required to travel the 160m mentioned
3.the time required to attain the speed of 30m/s
4.the distance moved from rest to the time time the train had a speed of 30m/s.
1
Expert's answer
2018-10-31T13:56:09-0400

The acceleration

a=(v_2^2-v_1^2)/(2s_1 )=(〖50〗^2-〖30〗^2)/(2×160)=5 m/s^2

The time required to travel the 160 m mentioned

t_1=(2s_1)/(v_1+v_2 )=(2×160)/(30+50)=4 s

The time required to attain the speed of 30 m/s

t=v_1/a=30/5=6 s

The distance moved from rest to the time the train had a speed of 30 m/s

s=(at^2)/2=(5×6^2)/2=90 m